Find a number k such that the following limit exist Also find the walue of the limit
Answer #1

( جے X X2 Le line 3x? + Kx +k+3 X-2 X² - x -2 lim 3x2+kx +k+3 = lim 3x2+kx+K+3 *32 (x-2)(x+1) x 2-2x + X-2 = lime (6) g(1) g